# Tight Algorithm for Connected Odd Cycle Transversal Parameterized by Clique-width

### 12 Feb 2024 v1

@@ 7 Conclusion
In this work, we have provided a single-sided error Monte Carlo algorithm for the Connected Odd Cycle Transversal problem with running time $\mathcal{O}^{*}(12^{\operatorname{cw}})$, presenting a new application of the counting representation approach by Bojikian and Kratsch [6]. We proved that the running time is tight assuming SETH. This closes the second open problem posed by Hegerfeld and Kratsch [21]. However, one more interesting benchmark problem is still open, namely the Feedback Vertex Set problem. As mentioned in their paper, solving this problem using the Cut&Count technique is challenging, since a direct application would require counting edges induced by partial solutions. Another approach would be to follow the approach of Bojikian and Kratsch [7]. The difficulty of such application lays in finding a low $\operatorname{GF}(2)$-rank representation of partial solutions.
As both the Cut&Count technique [13] and the rank-based approach [5] were used to solve connectivity problems more efficiently under different structural parameters, one would ask whether the approach of Bojikian and Kratsch [7] can also be adjusted to different settings, resulting in tight bounds for connectivity problems under different parameters, where there also exist a gap between the rank of a corresponding matrix, and the size of a largest triangular submatrix thereof.

### 26 Feb 2024 v2

@@ 7 Conclusion
In this work, we have provided a single-sided error Monte Carlo algorithm for the Connected Odd Cycle Transversal problem with running time $\mathcal{O}^{*}(12^{\operatorname{cw}})$, presenting a new application of the ”representative isolation” technique introduced by Bojikian and Kratsch [6]. We proved that the running time is tight assuming SETH. This closes the second open problem posed by Hegerfeld and Kratsch [21]. However, one more interesting benchmark problem is still open, namely the Feedback Vertex Set problem. As mentioned in their paper, solving this problem using the Cut&Count technique is challenging, since a direct application would require counting edges induced by partial solutions. Another approach would be to follow the approach of Bojikian and Kratsch [7]. The difficulty of such application lays in finding a low $\operatorname{GF}(2)$-rank representation of partial solutions.
Since both related techniques, namely the Cut&Count technique [13] and the rank-based approach [5], were used to solve connectivity problems more efficiently under different structural parameters, one would ask whether the approach of Bojikian and Kratsch [7] can also be adjusted to different settings, resulting in tight bounds for connectivity problems under different parameters, where there also exists a gap between the rank of a corresponding consistency matrix, and the size of a largest triangular submatrix thereof.
Acknowledgement.
The authors are grateful to Vera Chekan for her detailed comments on presentation, and to Falko Hegerfeld for several discussions on the lower bound presented here.